COMPLETE 2MM SPECTRAL LINE SURVEY (130-170 GHZ) OF SGR B2N, SGR B2OH, IRC +10 216, ORION (KL), ORION-S, W51M AND W3(IRS5)

\underline{ANTHONY J. REMIJAN}, DIANE P. LEIGH, BARRY E. TURNER, National Radio Astronomy Observatory, 520 Edgemont Road, Charlottesville, VA 22903; ANDREW J. MARKWICK-KEMPER, University of Manchester, Sackville Street, PO Box 88, Manchester M60 1QD, UK.

We report on a complete 2mm spectral line survey (130-170 GHz) taken with the NRAO^{1} 12m telescope between 1993 and 1995 toward the following sources: Sgr B2N, Sgr B2OH, IRC +10 216, Orion (KL), Orion-S, W51M and W3(IRS5).

Until very recently, this project was entirely the work of B.E. Turner. He wrote the original proposal, given in astro-ph document 0802.2273 (http://arxiv.org/abs/0802.2273) without changes or updates, and did all of the observing. Dr. Turner has fallen seriously ill and can no longer continue to work on the analysis of these data. The notes contained in the astro-ph document will also give further information about the project and important information for users of these data.

This is Barry’s Legacy Survey and as such, we felt it was essential to get the data out to the community for their use with the caveat that all the requisite citations and credit be given to Barry Turner and his collaborators. In this presentation, I will present the composite spectrum toward select sources and a brief analysis of the data toward Orion (KL) and IRC +10 216 including a comparison to corresponding survey with the IRAM 30m telescope. Any questions about the survey including obtaining the raw data should be sent to Anthony Remijan at aremijan@nrao.edu.

We fully expect that the users of these data give B. E. Turner and this work the appropriate citation and credit.

1The National Radio Astronomy Observatory is a facility of the National Science Foundation operated under cooperative agreement by Associated Universities, Inc.